The Magellanic Clouds Newsletter

⋆ ⋆ ⋆

An electronic exchange on Magellanic Clouds research

Edited by Bryan C. Dunne Co-Editors Eva K. Grebel and You-Hua Chu

                               mcnews@astro.uiuc.edu
                http://www.astro.uiuc.edu/projects/mcnews/MCNews.php

No. 77 July 14th, 2004

Letter From The Editor

Dear Colleagues,

The Magellanic Clouds Working Group is pleased to bring you issue 77 of the Magellanic Clouds Newsletter. In this issue, we present 6 submitted abstracts, an announcement for an upcoming meeting on science with large surveys, and a collection of recent astro-ph listings of potential interest to Magellanic Clouds researchers. As always, this month’s issue is available from the MC News Website in a variety of formats (PDF, HTML, PostScript, and L^{A}TEX).

Best Regards, Bryan Dunne Editor, MC News

Contents

6 Abstracts of Refereed Papers . . . . . . . . . . . . . 2 1 Meeting Announcement . . . . . . . . . . . . . . . . . . . 6 Recent astro-ph Listings . . . . . . . . . . . . . . . . . . . . 6

Upcoming Meeting Quicklist

Astrophysics in the Far Ultraviolet: Five Years of Discovery with FUSE . . . . . . . . . August 2–6

Astrophysics and Cosmology After Gamow - Theory and Observations . . . . . . . August 8–14

Diffuse Matter in the Galaxy: Observations Confront Theory . . . . . . . . . . . . . Aug 29–Sept 2

Science with LSST and Other Large Surveys . . Sept 20–22

Abstracts of Refereed Papers

Supernova Remnants in the Magellanic Clouds. IV. X-Ray Emission from the Largest SNR in the LMC

R. M. Williams (1), Y.-H. Chu (1), J. R. Dickel (1), R. A. Gruendl (1), R. Shelton (2), S. D. Points (3), R. C. Smith (3)

(1) Department of Astronomy, University of Illinois at Urbana-Champaign, Urbana, IL, USA (2) Deaprtment of Physics and Astronomy, University of Georgia, Athens, GA, USA

(3) Cerror-Telolo Interamerican Observatory, La Serena, Chile

We present the first X-ray detection of SNR 0450-70.9 the largest known supernova remnant (SNR) in the Large Magellanic Cloud. To study the physical conditions of this SNR, we have obtained XMM-Newton X-ray observations, optical images and high-dispersion spectra, and radio continuum maps. Optical images of SNR 0450-70.9 show a large, irregular elliptical shell with bright filaments along the eastern and western rims and within the shell interior. The interior filaments have higher [S II]/Halpha ratios and form an apparent inner shell morphology. The X-ray emission region is smaller than the full extent of the optical shell, with the brightest X-ray emission found within the small interior shell and on the western rim of the large shell. The expansion velocity of the small shell is ∼220 km/s, while the large shell is ∼120 km/s.

The radio image shows central brightening and a fairly flat radio spectral index over the SNR. However, no point X-ray or radio source corresponding to a pulsar is detected and the X-ray emission is predominantly thermal. Therefore, these phenomena can be most reasonably explained in terms of the advanced age of the large SNR. Using hydrodynamic models combined with a nonequilibrium ionization model for thermal X-ray emission, we derived a lower limit on the SNR age of about 45,000 yr, well into the later stages of SNR evolution. Despite this, the temperature and density derived from spectral fits to the X-ray emission indicate that the remnant is still overpressured, and thus that the development is largely driven by hot gas in the SNR interior. Comments: Accepted by ApJ WWW: http://arxiv.org/abs/astro-ph/0406295 e-mail: rosanina@astro.uiuc.edu

Three new X-ray pulsars detected in the Small Magellanic Cloud and the positions of two other known pulsars determined.

W. R. T. Edge,(1) M. J. Coe,(1) J. L. Galache,(1) V. A. McBride,(1) R.H.D. Corbet,(2 and 3) C.B. Markwardt,(2 and 4) S. Laycock (5)

(1) School of Physics and Astronomy, Southampton University

(2) NASA Goddard Space Flight Center

(3) Universities Space Research Association

(4) Department of Astronomy, University of Maryland (5) Harvard-Smithsonian Center for Astrophysics

Three new X-ray pulsars have been detected in the Small Magellanic Cloud (SMC) and the positions of two others have been determined, with archive Chandra data. A series of five observations of the SMC took place between May and October 2002. Analysis of these data has revealed three previously unknown X-ray pulsars at pulse periods of 34, 503 and 138 seconds. The position of pulsar XTE J0052-725, which was originally detected by RXTE on June 19 2002, was also accurately determined and a previously detected 7.78s RXTE pulsar was identified as the source SMC X-3.

Comments: Accepted by MNRAS WWW: http://arxiv.org/abs/astro-ph/0406580 e-mail: wrte@astro.soton.ac.uk

The young pulsar PSR B0540-69.3 and its synchrotron nebula in the optical and X-rays

N. I. Serafimovich (1,2), Yu. A. Shibanov (1), P. Lundqvist (2), J. Sollerman (2)

(1) Ioffe Physical Technical Institute, St. Petersburg, Russia

(2) Stockholm Observatory, AlbaNova Science Center, Department of Astronomy, Stockholm, Sweden

The young PSR B0540-69 (B0540) in the LMC is the only pulsar (except the Crab pulsar) for which a near-UV spectrum has been obtained. However, the absolute flux and spectral index of previously published HST/FOS data are significantly higher than suggested by broadband groundbased UBVRI photometry. Using our ESO/VLT/FORS1 spectral observations and HST/WFPC2 archival images we show that the old HST and new VLT spectral data are ¿50% contaminated by the Pulsar Wind Nebula (PWN) and that this is the reason for the above mentioned difference. We find that the broadband HST spectrum for the range 3300-8000 A is clearly nonthermal and has a negative spectral index of 1.07(+0.20/-0.19). This is different from the almost flat spectrum of the Crab pulsar. The PWN of B0540 shows a clear asymmetry of the surface brightness distribution along the major axis of the PWN torus-like structure with respect to the pulsar position, also seen in Chandra X-ray images. This can be linked to the asymmetry of the surrounding SN ejecta. We find no significant spectral index variation over the PWN. Using HST archival images we estimate the proper motion of B0540 to be 4.9+/-2.3 mas/yr, i.e. a transverse velocity of 1190+/-560 km/s along the southern jet of the PWN. This can make PSR B0540 the third pulsar with a proper motion aligned with the jet axis of its PWN, which poses constraints on pulsar kick models. We discuss the interstellar absorption toward B0540 including the contributions from the Milky Way, LMC and the supernova ejecta, and compare unabsorbed multiwavelength spectra of B0540 and the Crab pulsar, and their PWNs. Compared with the Crab, B0540 and its PWN show a significant depression in the optical range.

Comments: Accepted by A&A WWW: ftp://ftp.astro.su.se/pub/peter/0540/ e-mail: natalia@astro.su.se

A 2dF survey of the Small Magellanic Cloud

C. J. Evans (1), I. D. Howarth (2), M. J. Irwin (3), A. W. Burnley (2), T. J. Harries (4)

(1) Isaac Newton Group of Telescopes, La Palma

(2) Dept. of Physics & Astronomy, University College London

(3) Institute of Astronomy, University of Cambridge

(4) School of Physics, University of Exeter

We present a catalogue of new spectral types for hot, luminous stars in the Small Magellanic Cloud. The catalogue contains 4161 objects, giving an order of magnitude increase in the number of SMC stars with published spectroscopic classifications. The targets are primarily B- and A-type stars (2862 and 853 objects respectively), with 1 Wolf-Rayet, 139 O-type, and 306 FG stars, sampling the main sequence to ∼mid-B. The selection and classification criteria are described, and objects of particular interest are discussed, including UV-selected targets from the Ultraviolet Imaging Telescope (UIT) experiment, Be and B[e] stars, ‘anomalous A supergiants’, and composite-spectrum systems. We examine the incidence of Balmer-line emission, and the relationship between Hγ equivalent width and absolute magnitude for BA stars.

Comments: Accepted by MNRAS WWW: http://www.ing.iac.es/c̃je/2dF.html e-mail: cje@ing.iac.es

The Optical Gravitational Lensing Experiment. Small Amplitude Variable Red Giants in the Magellanic Clouds

I. Soszynski, A. Udalski, M. Kubiak, M. Szymanski, G. Pietrzynski, K. Zebrun, O. Szewczyk, L. Wyrzykowski

We present analysis of the large sample of variable red giants from the Large and Small Magellanic Clouds detected during the second phase of the Optical Gravitational Lensing Experiment (OGLE-II) and supplemented with OGLE-III photometry. Comparing pulsation properties of detected objects we find that they constitute two groups with clearly distinct features. In this paper we analyze in detail small amplitude variable red giants (about 15400 and 3000 objects in the LMC and SMC, respectively). The vast majority of these objects are multi-periodic. At least 30% of them exhibit two modes closely spaced in the power spectrum, what likely indicates non-radial oscillations. About 50% exhibit additional so called Long Secondary Period. To distinguish between AGB and RGB red giants we compare PL diagrams of multi-periodic red giants located above and below the tip of the Red Giant Branch (TRGB). The giants above the TRGB form four parallel ridges in the PL diagram. Among much more numerous sample of giants below the TRGB we find objects located on the low luminosity extensions of these ridges, but most of the stars are located on the ridges slightly shifted in logP. We interpret the former as the second ascent AGB red giants and the latter as the first ascent RGB objects. Thus, we empirically show that the pulsating red giants fainter than the TRGB are a mixture of RGB and AGB giants. Finally, we compare the Petersen diagrams of the LMC, SMC and Galactic bulge variable red giants and find that they are basically identical indicating that the variable red giants in all these different stellar environments share similar pulsation properties.

Comments: Published in Acta Astronomica 54 (2004) 129 WWW: http://arxiv.org/abs/astro-ph/0407057 e-mail: soszynsk@astrouw.edu.pl

The Low- and Intermediate-Mass Stellar Population in the Small Magellanic Cloud: The Central Stars of Planetary Nebulae

Eva Villaver(1,2), Letizia Stanghellini(3) and Richard A. Shaw(3)

(1)Space Telescope Science Institute (2)Affiliated with the Hubble Space Telescope Division of the European Space Agency (3)National Optical Astronomy Observatory

We present a study on the central stars (CSs) of Planetary Nebulae (PNe) observed in the Small Magellanic Cloud (SMC) with the Space Telescope Imaging Spectrograph instrument on-board the HST. The stellar magnitudes have been measured using broad-band photometry, and Zanstra analysis of the nebulae provided the stellar temperatures. From the location of the CSs on the HR diagram, and by comparing the observed CSs with current models of stellar evolution, we infer the CSs masses. We examine closely the possibility of light contamination in the bandpass from an unrecognized stellar companion, and we establish strong constraints on the existence and nature of any binary companion. We find an average mass of 0.63 Msun, which is similar to the mass obtained for a sample of CSs in the LMC (0.65 Msun). However, the SMC and LMC CS mass distributions differ slightly, the SMC sample lacking an intermediate-mass stellar population (0.65 to 0.75 Msun). We discuss the significance and possible reasons for the difference between the two mass distributions. In particular, we consider the differences in the star formation history between the clouds and the mass-loss rate dependence on metallicity

Comments: To appear in the October 20 issue of the Astrophysical Journal WWW: http://arxiv.org/abs/astro-ph/0407033 e-mail: villaver@stsci.edu

Meeting Announcements

Science with LSST and Other Large Surveys: Community Access and Utilization of Archives of the Future

September 20–22, 2004 University of Washington, Seattle, USA

The Large Synoptic Survey Telescope (LSST) is in its preliminary design phase. The functional and performance requirements of the telescope are being developed now from a set of scientific experiments that it will carry out starting as early as 2012, as recommended by the most recent Decadal Survey of astronomy. LSST will produce a rich and extensive public archive of time-resolved astronomical data. The LSST archive will be of unprecedented size, and will be immediately available to the general community and the public. This meeting in September 2004 is an opportunity to learn about the current scientific thinking connected with LSST, to explore and discuss additional areas for which the LSST dataset might provide valuable information, and to contribute to the definition of the infrastructure and tools that will ensure the legacy of LSST. These discussions will be carried out in the context of other large, optical/infrared survey projects that are in various stages of planning or implementation.

There will be invited and contributed presentations and significant time for discussion. For more detailed program information, registration or to submit a contributed talk or poster please visit http://www.lsst.org/Meetings/CommAccess/.

Deadlines ∘ Registration deadline: August 31st, 2004

For more information, contact the meeting organizers at: WWW: http://www.lsst.org/Meetings/CommAccess/ E-mail: graybeal@astro.washington.edu or kcook@guen.ucllnl.org

Recent astro-ph Listings

astro-ph/0407258 Title: A Be star with a low nitrogen abundance in the SMC cluster NGC330 Authors: D. J. Lennon (1), J.-K. Lee (2), P. L. Dufton (2), R. S. I. Ryans (2) ((1) Isaac Newton Group,

(2) Queen’s University Belfast)

Comments: 7 pages, 4 figures. Submitted to A&A astro-ph/0407226

Title: The young pulsar PSR B0540-69.3 and its synchrotron nebula in the optical and X-rays Authors: N. I. Serafimovich, Yu. A. Shibanov, P. Lundqvist, J. Sollerman Comments: 21 pages, 15 figures, A&A accepted (paper can be also be loaded from this ftp URL)

astro-ph/0407210 Title: Microlensing towards the LMC Authors: Philippe Jetzer Comments: 10 pages, 4 figure, Proceedings of the conference ”Thinking, Observing and Mining the Universe (September 2003, Sorrento) astro-ph/0407209

Title: Microlensing towards LMC and M31 Authors: Philippe Jetzer, Sebastiano Calchi Novati Comments: 18 pages, 8 figures. to appear in the proceedings of the Marcel Grossmann Meeeting X

astro-ph/0407115 Title: Origin of the Soft Excess in X-ray Pulsars Authors: Ryan C. Hickox (1), Ramesh Narayan (1), Timothy R. Kallman (2) ((1) Harvard-Smithsonian Center for Astrophysics, (2) NASA Goddard Space Flight Center) Comments: 17 pages, 9 figures, accepted for publication in ApJ

astro-ph/0407057 Title: The Optical Gravitational Lensing Experiment. Small Amplitude Variable Red Giants in the Magellanic Clouds Authors: I. Soszynski, A. Udalski, M. Kubiak, M. Szymanski, G. Pietrzynski, K. Zebrun, O. Szewczyk, L. Wyrzykowski Comments: 23 pages, 13 figures included in GIF format Journal-ref: Acta Astron. 54 (2004) 129

astro-ph/0407033 Title: The Low- and Intermediate-Mass Stellar Population in the Small Magellanic Cloud: The Central Stars of Planetary Nebulae Authors: Eva Villaver (STScI/ESA), Letizia Stanghellini (NOAO), Richard A. Shaw (NOAO) Comments: 30 pages, 6 figures, 5 tables. To be published in ApJ (October 20)

astro-ph/0407006 Title: The Magellanic Clouds Photometric Survey: The Large Magellanic Cloud Stellar Catalog and Extinction Map Authors: Dennis Zaritsky (U. Arizona), Jason Harris (U. Arizona), Ian B. Thompson (OCIW), Eva K. Grebel (U. Basel) Comments: 7 pages (emulateapj), accepted for publication in the Astronomical Journal

astro-ph/0406580 Title: Three new X-ray pulsars detected in the Small Magellanic Cloud and the positions of two other known pulsars determined Authors: W. R. T. Edge, M. J. Coe, J. L. Galache, V. A. McBride, R.H.D. Corbet, C.B. Markwardt, S. Laycock Comments: Accepted by MNRAS

astro-ph/0406475 Title: Cluster formation versus star formation rates around six regions in the LMC Authors: Annapurni Subramaniam Comments: to appear in A&A, 12 pages including 7 figures

astro-ph/0406443 Title: Explaining the mysterious age gap of globular clusters in the Large Magellanic Cloud Authors: K. Bekki, W. J. Couch, M. A. Beasley, D. A. Forbes, M. Chiba, G. S. Da Costa Comments: 14 pages, 4 figures (f2.copy.jpg for high-res color figure), accepted by ApJL

astro-ph/0406420 Title: Long Term Behavior of X-ray Pulsars in the Small Magellanic Cloud Authors: S. Laycock (CfA), R. H. D. Corbet (GSFC, USRA), M. J. Coe (U. Southampton), F. E. Marshall (GSFC), C. Markwardt (GSFC, U. Maryland), J. Lochner (GSFC) Comments: 31 pages, submitted to ApJ

astro-ph/0406409 Title: A 2dF survey of the Small Magellanic Cloud

Magellanic Clouds Newsletter No. 77  7

Authors: C. J. Evans, I. D. Howarth, M. J. Irwin, A. W. Burnley, T. J. Harries Comments: 27 pages, 17 figures (reduced size). Accepted for publication in MNRAS. Both a postscript version with normal res. figs and the full catalogue is available from this http URL

astro-ph/0406295 Title: Supernova Remnants in the Magellanic Clouds. IV. X-Ray Emission from the Largest SNR in the LMC Authors: R. M. Williams, Y.-H. Chu, J. R. Dickel, R. A. Gruendl, R. Shelton, S. D. Points, R. C. Smith Comments: Accepted for publication in The Astrophysical Journal

astro-ph/0406197 Title: AST/RO Observations of CO J=4-3 Emission from the N44 Complex in the Large Magellanic Cloud Authors: Sungeun Kim, Wilfred Walsh, Kecheng Xiao Comments: 11 pages, 5 figures, submitted to ApJ

astro-ph/0406170 Title: IRS observations of the LMC planetary nebula SMP83 Authors: J. Bernard-Salas (1), J.R. Houck (1), P.W. Morris (2), G.C. Sloan (1), S.R. Pottasch (3), D.J. Barry (1), ((1) Cornell University, (2) Caltech/SSC, (3) Kapteyn Astronomical Institute) Comments: Accepted in ApJSup (Spitzer Special Issue), 4 pages, 2 figures

astro-ph/0406040 Title: Spitzer Space Telescope Observations of the Aftermath of Microlensing Event MACHO-LMC-5 Authors: Hien Trong Nguyen, Nitya Kallivayalil, Michael Werner, Charles Alcock, Brian Patten, Daniel Stern Comments: Accepted for publication, ApJS Spitzer Special Issue

astro-ph/0406015 Title: XMM-Newton Observations of the Be/X-ray transient A0538-66 in quiescence Authors: P. Kretschmar (MPE/ISDC), J. Wilms (Univ. Warwick), R. Staubert (IAA Tuebingen), I. Kreykenbohm (IAAT/ISDC), W. A. Heindl (CASS/UCSD) Comments: Accepted for proceedings of 5th INTEGRAL Workshop

astro-ph/0405555 Title: Gravity, Turbulence, and Star Formation Authors: Bruce G. Elmegreen (IBM T.J. Watson Research Center) Comments: 6 pages, 1 figure, In Star Formation in the Interstellar Medium: A 60th birthday celebration for David Hollenbach, Chris McKee, and Frank Shu, Ed. F. Adams, D. Johnstone, D. Lin and E. Ostriker, Astronomical Society of the Pacific (PASP Conference series), in press, from a conference at Lake Tahoe, CA, June 29-July 3 2003

astro-ph/0405551 Title: EROS 2 photometry of probable R Coronae Borealis stars in the Small Magellanic Cloud Authors: P. Tisserand, J.B. Marquette, J.P. Beaulieu, P. de Laverny, et al. (The EROS collaboration) Comments: 8 pages, 9 figures, accepted for publication in A&A. Uses aa.cls and natbib.sty files

Magellanic Clouds Newsletter No. 77  8